# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v1

@@ Abstract
In this paper, we study the well-posedness of Fractional Rough Burgers equation driven by space-time noise in $H^{s}(\mathbb{T})$ space. For the higher dissipation $\gamma\in(\frac{4}{3},2]$, we establish local well-posedness. Global well-posedness is further obtained when $\gamma$ is restricted to the interval $(\frac{3}{2},2]$. For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, we use the regularity analysis derivation the para-controlled solution.
@@ 1 Introduction
$\partial_{t}u-\Lambda^{\gamma}u=\frac{1}{2}\partial_{x}(u^{2})+\xi,\quad u(0,x)=u_{0}.$ (1.1)
Here $\mathbb{T}$ is some toru satisfied $\int_{\mathbb{T}}e^{ikx}=0$ for any $k\in\mathbb{Z}/\{0\}$(for example $\mathbb{R}/2\pi\mathbb{Z}$) and $\Lambda^{\gamma}$ is a Fourier multiplier satisfied $\mathcal{F}(\Lambda^{\gamma}f)=-|\xi|^{\gamma}\mathcal{F}f$. $\xi$ is a centered Gaussian space-time random distribution with covariance
@@
Equation (1.1) is a classical model in SPDE driving by additional noise. Similar research topics have attracted extensive study in recent years, building upon the foundational work of Martin Hairer in [Hai13, Hai14] and Gubinelli, Perkowski in [GIP15, GP17]. In this paper, we primarily investigate how the degree of dissipation affects the equation and solutions. For the high dissipation which $\frac{5}{3}<\gamma\leq 2$, the Fractional Rough Burgers equation with nonlinearity $F(u)\partial_{x}u$ replacing $\frac{1}{2}\partial_{x}(u^{2})$ has been studied in [GIP15]. Additionally, they build the global well-posedness in $\mathcal{C}^{s}$ for $\frac{1}{3}<s<\frac{\gamma}{2}-\frac{1}{2}$.
Note that the performance of lower dissipation $\gamma<\frac{5}{3}$ in $H^{s}$ space is unknown, where Sobolev spaces $H^{s}$ are of significant mathematical and physical relevance in the study Burgers equation and Degasperis-Procesi(DP) equation. For more details, one can be refer to [Yin03, Yin04, GL11, GL15]. In this paper, we will consider equation (1.1) with lower dissipation $\frac{4}{3}<\gamma\leq 2$ for $u_{0}\in H^{\frac{1}{2}+}$ and $\frac{5}{4}<\gamma\leq\frac{4}{3}$ for $u_{0}\in H^{\frac{1}{2}+}(\mathbb{T})\cap\mathcal{C}^{\frac{1}{2}+}(\mathbb{T})=W^{\frac{1}{2}+}(\mathbb{T})$.
It is also need to particularly noteworthy that the first author and co-authors has previously studied the randomized initial data problems for both the DP equation and Burgers equation
@@
The primary objective of this paper is to establish fundamental results for the FRB equation in Sobolev spaces $H^{s}$ and extend analogous conclusions to the DP equations. We will address the potential difficulties associated with these problems, as well as the methods we will employ, by organizing the discussion into three main sections.
@@ 1.1 Probabilistic setting
To have a better understanding of the problem, we first study (1.1) by writing it as mild form
$u=P(t)u_{0}+\frac{1}{2}\int_{0}^{t}P(t-s)\partial_{x}(u^{2})\mathrm{d}s+X(t).$ (1.4)
@@
Note that $X(t)\in C_{T}\mathcal{C}^{\alpha}(\mathbb{T})$ a.s for any $\alpha<\frac{\gamma}{2}-\frac{1}{2}$, which is applied to the study of equations describing thermal dissipative structures such as KPZ equation [Hai13] and Burgers equation [GIP15, GP17] to establish well-posedness. This also applies to Sobolev spaces $H^{s}$ (See Lemma 2.7) and keep in mind that $X\in C_{T}W^{\alpha}(\mathbb{T})$ for $\alpha<\frac{\gamma}{2}-\frac{1}{2}$.
Note that the range of $\alpha$ is determined by dissipation $\gamma$, our paper gives two works in different range. One is higher regularity $\alpha\in(\frac{1}{6},\frac{1}{2})$, in such case, we can directly establish the well-posedness in $H^{\alpha}(\mathbb{T})$. One is lower dissipation $\alpha\in(\frac{1}{8},\frac{1}{6})$(See Remark 3.1), in such case we need to work in $W^{\alpha}(\mathbb{T})$. A common feature is that we have isolated a higher-regularity term (More precisely $H^{\frac{1}{2}+}(\mathbb{T})$) in the solution. This process is made possible in lower disspation $\gamma\in(\frac{5}{4},\frac{4}{3}]$ with the assistance of para-controlled method which is a powerful method to study rough or singular SPDE in [GIP15, GP17, GKO23, CC18].
@@
$u_{t}-\Lambda^{\gamma}u=\frac{1}{2}\partial_{x}(u^{2})+|D|^{\beta}\xi,\quad u(0,x)=u_{0}$ (1.6)
@@ 1.2 Regularity analysis
$\partial_{t}v-\Lambda^{\gamma}v=\frac{1}{2}\partial_{x}(v^{2})+f,\quad v(0,x)=v_{0},$ (1.8)
where $f$ may depend on $u$. It takes the different structure of solution in two different range $\gamma\in(\frac{4}{3},2]$ and $\gamma\in(\frac{5}{4},\frac{4}{3}]$.
More precisely, let $f(v)=\partial_{x}(vX)+\frac{1}{2}\partial_{x}(X^{2})$, where $X$ satisfied (1.5). Then $f(v)$ gains at most $\alpha$-order Sobolev regularity when $v$ belongs to $H^{\frac{1}{2}+}$. It derive the relation between $\alpha$ and $\gamma$ which satisfied $\alpha+\gamma\geq\frac{3}{2}$, since $v$ can gain $\alpha+\gamma-1$ at most. For the lower $\alpha$ and $\gamma$, we use the Little-wood Paley decomposition in [BCD11] which rewrite $vX$ and $X^{2}$ as
@@
It takes the higher regularity $T_{X}v$, $R(v,X)$ and $R(X,X)$ gains since $X$ has $\alpha$-order Höler regularity. The only difficult for the case of smaller $\alpha$ is $T_{v}X$ and $2T_{X}X$ which can be countered by para-controlled solution(See more details in Proposition 4.1), and it derives the relation $2\alpha-1+\gamma\geq\frac{3}{2}$.
@@ 1.3 Shallow water equation
Next, we introduce the Degasperis-Procesi(DP) equation, which is one type of shallow water equation, has following form
@@
Then we can define the fluid velocity $u(t,x):=g\ast m:=(1-\partial_{x}^{2})m$ and rewrite the DP equation as
@@
We provide following citation [DP99, DHH02] to facilitate a deeper understanding of the physical background for this class of equations. In particular, a systematic discussion can be found in [AG10]. For the mathematical properties of this equation, we refer readers to the following literature. Of particular note, the local well-posedness to (1.10) with initial data $H^{s}$, $s\geq\frac{3}{2}$ was proved in [LY06], and also the precise blow-up scenario and a blow-up result were derived in [LY07].
@@
$u_{t}+\Lambda^{\gamma}u+\frac{1}{2}\partial_{x}(u^{2})+\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})=\xi,\quad u(0,x)=u_{0}(x).$ (1.12)
as an application of our result for Rough Burgers equation.
@@ 1.4 Main result
Let $u_{0}(x)\in H^{\frac{1}{2}+\delta}(\mathbb{T})$, for some small $\delta>0$, and $\gamma\in(\frac{4}{3},2]$, then there exists a postive $T^{*}$ and a solution $u(t,x)$ of equation (1.1) satisfied
@@
where $v(t,x)\in C_{T^{*}}H^{\frac{1}{2}+\delta}(\mathbb{T})$ a.s. and $X(t,x)$ is satisfied (1.5) with $X(t,x)\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ for $\alpha\in(\frac{1}{6},\frac{\gamma}{2}-\frac{1}{2})$, a.s. Moreover we have $u\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s.
@@
’a.s.’ represent almost surely. We say event $A$ is almost surely if and only if $\mathbb{P}(A)=1$, where $\mathbb{P}$ represent the probability associated with $(\Omega,\mathcal{F},\mathbb{P})$. We will omit this description in most situation, except we need distinct such meaning.
@@
where $G\in C_{b}^{3}(\mathbb{R}^{d},L(\mathbb{R}^{d},\mathbb{R}^{d}))$. The main difference between our result and [GIP15] lies in the fact that our solution can back to the original equation (1.1). Since $G(u)\partial_{x}u$ is ill-defined when $\alpha\in(\frac{1}{3},\frac{\gamma}{2}-\frac{1}{2})$, they actually study the “generalized" solution, which is the limit in probability sense of $u^{\epsilon}$ satisfied following mollified equation
@@
Here $\xi^{\epsilon}=\epsilon^{-1}\psi(\epsilon\cdot)\ast\epsilon$ for some $\psi\in\mathcal{S}$ with $\int\psi\mathrm{d}t=1$. Similar approaches are frequently used for some singular SPDEs in [CC18, Hai13, GP17].
@@
For $\gamma>\frac{3}{2}$, Rough Burgers equation (1.1) has a global weak solution $u=v+X$ in $C([0,T];L^{2}(\mathbb{T}))\cap L^{2}([0,T];H^{\frac{\gamma}{2}}(\mathbb{T}))$ for any $T\geq 0$ . Moreover, for $\gamma>\frac{5}{3}$, such solution is the global strong solution of (1.1).
@@
For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, let $\mathcal{Q}$ be the linear evolution of $\partial_{x}X$ such as
@@
We let $(u,u^{\prime}u^{\sharp})$ be a triple satisfied
@@
Let $u_{0}\in\mathcal{C}^{\frac{1}{2}+\delta}\cap H^{\frac{1}{2}+\delta}=W^{\frac{1}{2}+\delta}$ for some $\delta>0$, $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Define $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then there exists a positive $T^{*}$ and a unique triple $(u,u^{\prime},u^{\sharp})\in C_{T}W^{\alpha}\times C_{T}W^{\frac{1}{8}+\delta}\times C_{T}W^{\frac{1}{2}+\delta}$ satisfied
$u^{\prime}=$ $X+2u^{\prime}\prec\!\!\!\prec u^{\sharp},$ (1.17)
@@
where $X(t)$ is satisfied (1.5). Moreover, $u$ is a solution of (1.1).
@@
This is a term that will appear in the equation for $u^{\sharp}$. Our strategy is to use a specific $u^{\prime}$ with $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})$ to cancel out this term to make sure that $u^{\sharp}$ can gains higher regularity (See Proposition 4.1), we will show it in Section 4.
@@
We can extend the range of $\gamma\in(1,2]$, which means the $\alpha\in(0,\frac{1}{2})$. We omit these proofs because in this paper, since we primarily aim to demonstrate the effectiveness of the paracontrolled method in enhancing the regularity of solutions.
@@ 1.5 Organization of the paper
This paper is organized as follows. In Section 2, we introduce some notations and basic tools. In Section 3, we prove Theorem 1.1 and Theorem 1.2 for FRB equation and Remark 1.3 for Rough DP equation. In the Section 4, we prove Theorem 1.3 for para-controlled solution. In the Section 5, we give a conjecture of the range of $\gamma$ and give some future directions.

### v2

@@ Abstract
In this paper, we study the well-posedness of Fractional Rough Burgers equation driven by space-time white noise in $H^{s}(\mathbb{T})$ space. For the higher dissipation $\gamma\in(\frac{4}{3},2]$, we establish local well-posedness. Global well-posedness is further obtained when $\gamma$ is restricted to the interval $(\frac{3}{2},2]$. For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, we use the regularity analysis derivation the para-controlled solution.
@@ 1 Introduction
$\partial_{t}u-\Lambda^{\gamma}u=\partial_{x}(u^{2})+\xi,\quad u(0,x)=u_{0}.$ (1.1)
Here $\mathbb{T}$ is some torus satisfied $\int_{\mathbb{T}}e^{ikx}=0$ for any $k\in\mathbb{Z}/\{0\}$(for example $\mathbb{R}/2\pi\mathbb{Z}$) and $\Lambda^{\gamma}$ is a Fourier multiplier satisfied $\mathcal{F}(\Lambda^{\gamma}f)=-|\xi|^{\gamma}\mathcal{F}f$. $\xi$ is a centered Gaussian space-time random distribution with covariance
@@
Equation (1.1) is a classical stochastic partial differential equation driven by additive noise. This class of problems has attracted considerable attention in recent years, largely motivated by the pioneering work on regularity structures by Hairer [Hai13, Hai14] and on paracontrolled calculus by Gubinelli and Perkowski [GIP15, GP17].
The central objective of this paper is to understand how the dissipation parameter $\gamma$ affects the well-posedness and regularity of solutions for the specific quadratic nonlinearity $\partial_{x}(u^{2})$. This nonlinearity is of particular interest due to its close connection to the KPZ equation; see, for example, [Cor12]. In contrast, previous work such as [GIP15] studied the fractional stochastic Burgers equation with a general nonlinearity $F(u)\partial_{x}u$, establishing well-posedness in $\mathcal{C}^{s}$ for $s\in(\frac{1}{3},\frac{\gamma}{2}-\frac{1}{2})$ in the regime $\gamma\in(\frac{5}{3},2]$.
The behavior of equation (1.1) with lower dissipation $\gamma<\frac{5}{3}$ in the Sobolev spaces $H^{s}$ remains largely unexplored. These spaces are of significant mathematical and physical relevance in the study of related equations such as the Burgers and Degasperis–Procesi equations (see, e.g., [Yin03, Yin04, GL11, GL15]). In this paper, we consider equation (1.1) under two regimes of lower dissipation:
  • For $\frac{4}{3}<\gamma\leq 2$, we assume initial data $u_{0}\in H^{\frac{1}{2}+}(\mathbb{T})$.
  • For $\frac{5}{4}<\gamma\leq\frac{4}{3}$, we work with initial data in the intersection $u_{0}\in H^{\frac{1}{2}+}(\mathbb{T})\cap\mathcal{C}^{\frac{1}{2}+}(\mathbb{T})=W^{\frac{1}{2}+}(\mathbb{T})$.
It is also need to particularly noteworthy that the first author and co-authors have previously studied the randomized initial data problems for both the DP equation and Burgers equation
@@
The primary objection of this paper is to establish fundamental results for the FRB equation in Sobolev spaces $H^{s}$ and extend analogous conclusions to the DP equations. We will address the potential difficulties associated with these problems, as well as the methods we will employ, by organizing the discussion into three main sections.
@@ 1.1 Probabilistic setting
The fact that $\xi\in\mathcal{C}^{-\frac{1}{2}-}(\mathbb{R}_{+}\times\mathbb{T})$ introduces significant analytical difficulties for direct approaches. One strategy is to write equation as mild form
$u=P(t)u_{0}+\int_{0}^{t}P(t-s)\partial_{x}(u^{2})\mathrm{d}s+X(t).$ (1.4)
@@
Noting that $X(t)\in C_{T}\mathcal{C}^{\alpha}(\mathbb{T})$ holds almost surely for any $\alpha<\frac{\gamma}{2}-\frac{1}{2}$, Such regularity has played a key role in establishing well-posedness for equations modeling thermally driven dissipative structures, notably the KPZ equation [Hai13] and the Burgers equation [GIP15, GP17]. In this paper, we extend analogous regularity results to the Sobolev spaces $H^{s}$ and apply them to prove well-posedness for the class of equations considered.
Since the admissible range of $\alpha$ depends on the dissipation parameter $\gamma$, this paper presents two regimes with distinct analytical treatments.
  • Higher regularity regime ($\alpha\in(\frac{1}{6},\frac{1}{2})$): in this case, well-posedness can be established directly in the Sobolev space $H^{\alpha}(\mathbb{T})$.
  • Lower dissipation regime ($\alpha\in(\frac{1}{8},\frac{1}{6})$, see Remark 3.1): here the analysis requires working in $W^{\alpha}(\mathbb{T})$.
A common feature in both regimes is the isolation of a higher-regularity component in the solution—more precisely, a term belonging to $H^{\frac{1}{2}+}(\mathbb{T})$. This decomposition is made possible for lower dissipation levels $\gamma\in(\frac{5}{4},\frac{4}{3}]$ by means of the paracontrolled calculus, a powerful tool introduced for studying rough and singular stochastic partial differential equations [GIP15, GP17, GKO23, CC18].
@@
$u_{t}-\Lambda^{\gamma}u=\partial_{x}(u^{2})+|D|^{\beta}\xi,\quad u(0,x)=u_{0}.$ (1.6)
@@ 1.2 Regularity analysis
$\partial_{t}v-\Lambda^{\gamma}v=\partial_{x}(v^{2})+f,\quad v(0,x)=v_{0},$ (1.8)
where $f$ may depend on $u$. It takes the different structure of solution in two different ranges $\gamma\in(\frac{4}{3},2]$ and $\gamma\in(\frac{5}{4},\frac{4}{3}]$.
More precisely, let $f(v)=2\partial_{x}(vX)+\partial_{x}(X^{2})$, where $X$ satisfied (1.5). Then $f(v)$ gains at most $\alpha$-order Sobolev regularity when $v$ belongs to $H^{\frac{1}{2}+}$. It derive the relation between $\alpha$ and $\gamma$ which satisfied $\alpha+\gamma\geq\frac{3}{2}$, since $v$ can gain $\alpha+\gamma-1$ at most.
For the lower $\alpha$ and $\gamma$, our objective is to similarly isolate a higher-regularity remainder term ($H^{\frac{1}{2}+}$) in the solution’s structure. As the method used in the highly dissipative regime is no longer applicable here, a more refined analysis is required. Using the Littlewood–Paley decomposition (see [BCD11]), we rewrite $vX$ and $X^{2}$ as
@@
The terms $T_{X}v$, $R(v,X)$ and $R(X,X)$ inherit higher regularity due to the fact that $X$ possesses $\alpha$-order Höler regularity. The main difficulty in the small-$\alpha$ case lies in handling the remaining terms $T_{v}X$ and $2T_{X}X$. We address this difficulty using the paracontrolled method (see Proposition 4.1 for details), which ultimately yields the condition $2\alpha-1+\gamma\geq\frac{3}{2}$.
@@ 1.3 Shallow water equation
Next, we introduce the Degasperis-Procesi(DP) equation, which is one type of shallow water equation, has the following form
@@
Then we can define the fluid velocity $u(t,x):=g\ast m:=(1-\partial_{x}^{2})^{-1}m$ and rewrite the DP equation as
@@
The following citations [DP99, DHH02] provide physical background for this class of equations; a more systematic discussion can be found in [AG10]. Regarding the mathematical theory, we note in particular that local well-posedness of (1.10) for initial data in $H^{s}$ with $s>\frac{3}{2}$ was established in [LY06], together with a precise blow-up criterion and blow-up results in [LY07].
@@
$u_{t}-\Lambda^{\gamma}u+\frac{1}{2}\partial_{x}(u^{2})+\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})=\xi,\quad u(0,x)=u_{0}(x),$ (1.12)
as an application of our result for Rough Burgers equation. Such method is also suitable for other type of shallow water equation with fractional dissipation.
@@ 1.4 Main result
Let $u_{0}(x)\in H^{\frac{1}{2}+\delta}(\mathbb{T})$ for $0<\delta<\frac{3\gamma}{8}-\frac{1}{2}$, $\gamma\in(\frac{4}{3},2]$, then there exists a postive $T^{*}$ and a solution $u(t,x)$ of equation (1.1) satisfying
@@
where $v(t,x)\in C_{T^{*}}H^{\frac{1}{2}+\delta}(\mathbb{T})$ a.s. and $X(t,x)$ is satisfied (1.5) with $X(t,x)\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s for $\alpha\in(\frac{1}{6},\frac{\gamma}{2}-\frac{1}{2})$. Moreover we have $u\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s..
@@
‘a.s.’ represent almost surely. We say event $A$ is almost surely if and only if $\mathbb{P}(A)=1$, where $\mathbb{P}$ represent the probability associated with $(\Omega,\mathcal{F},\mathbb{P})$. We will omit this description in most situation, except we need distinct such meaning.
@@
where $G\in C_{b}^{3}(\mathbb{R}^{d},L(\mathbb{R}^{d},\mathbb{R}^{d}))$. The main difference between our result and [GIP15] lies in the fact that our solution can back to the original equation (1.1). Since $G(u)\partial_{x}u$ is ill-defined when $\alpha\in(\frac{1}{3},\frac{\gamma}{2}-\frac{1}{2})$, they actually study the “generalized" solution, which is the limit in probability sense of $u^{\epsilon}$ satisfying the following mollified equation
@@
Here $\xi^{\epsilon}=\epsilon^{-1}\psi(\epsilon\cdot)\ast\xi$ for some $\psi\in\mathcal{S}$ with $\int\psi\mathrm{d}t=1$. Similar approaches are frequently used for some singular SPDEs in [CC18, Hai13, GP17].
@@
For $\gamma\in(\frac{3}{2},2]$ and $u_{0}\in H^{\frac{1}{2}+\delta}$ where $\delta$ satisfies Theorem 1.1, Rough Burgers equation (1.1) has a global weak solution $u=v+X$ where $X$ satisfies (1.5) and $v$ is belongs to $C([0,T];L^{2}(\mathbb{T}))\cap L^{2}([0,T];H^{\frac{\gamma}{2}}(\mathbb{T}))$ for any $T\geq 0$ . Moreover, for $\gamma>\frac{5}{3}$, such solution is the global strong solution of (1.1).
@@
For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, let $\mathcal{Q}$ be the linear evolution of $\partial_{x}X$ such that
@@
Let $(u,u^{\prime},u^{\sharp})$ be a triple satisfying
@@
Let $u_{0}\in\mathcal{C}^{\frac{1}{2}+\delta}\cap H^{\frac{1}{2}+\delta}=W^{\frac{1}{2}+\delta}$ for $0<\delta<\frac{\gamma}{2}-\frac{5}{8}$, $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Define $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then there exists a positive $T^{*}$ and a unique triple $(u,u^{\prime},u^{\sharp})\in C_{T}W^{\alpha}\times C_{T}W^{\frac{1}{8}+\delta}\times C_{T}W^{\frac{1}{2}+\delta}$ satisfied
$u^{\prime}=$ $2X+2u^{\prime}\prec\!\!\!\prec\mathcal{Q}+2u^{\sharp},$ (1.17)
@@
where $X(t)$ is satisfied (1.5) with $X(t,x)\in C_{T^{*}}W^{\alpha}(\mathbb{T})$ a.s for $\alpha\in(\frac{1}{8},\frac{\gamma}{2}-\frac{1}{2})$. Moreover, $u$ is a solution of (1.1).
@@
This term that will appear in the equation for $u^{\sharp}$. Our strategy is to use a specific $u^{\prime}$ with $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})$ to cancel out this term to make sure that $u^{\sharp}$ can gains higher regularity (See Proposition 4.1), we will show it in Section 4.
@@
We can extend the range to $\gamma\in(1,2]$, which means $\alpha\in(0,\frac{1}{2})$. We omit these proofs because in this paper, since we primarily aim to demonstrate the effectiveness of the paracontrolled method in enhancing the regularity of solutions.
@@ 1.5 Organization of the paper
This paper is organized as follows. In Section 2, we introduce some notations and basic tools. In Section 3, we prove Theorem 1.1 and Theorem 1.2 for the fractional stochastic Burgers equation, and we also prove the result in Remark 1.3 for the rough Degasperis–Procesi equation. In Section 4, we prove Theorem 1.3 for the paracontrolled solution. In Section 5, we provide a conjecture on the range of $\gamma$ in Sobolev space and discuss some future research directions.